# Stealing Trust: Unraveling Blind Message Attacks in Web3 Authentication

### v1

@@ Abstract.
As the field of Web3 continues its rapid expansion, the security of Web3 authentication, often the gateway to various Web3 applications, becomes increasingly crucial. Despite its widespread use as a login method by numerous Web3 applications, the security risks of Web3 authentication have not received much attention. This paper investigates the vulnerabilities in the Web3 authentication process and proposes a new type of attacks, dubbed blind message attacks. In blind message attacks, attackers trick users into blindly signing messages from target applications by exploiting users’ inability to verify the source of messages, thereby achieving unauthorized access to the target application. We have developed Web3AuthChecker, a dynamic detection tool that interacts with Web3 authentication-related APIs to identify vulnerabilities. Our evaluation of real-world Web3 applications shows that a staggering 75.8% (22/29) of Web3 authentication deployments are at risk of blind message attacks. In response to this alarming situation, we implemented Web3AuthGuard on the open-source wallet MetaMask to alert users of potential attacks. Our evaluation results show that Web3AuthGuard can successfully raise alerts in 80% of the tested Web3 authentications. We have responsibly reported our findings to vulnerable websites and have been assigned two CVE IDs.
@@
copyright: acmcopyrightconference: 31st ACM Conference on Computer and Communications Security; 2024journalyear: ;$c$s: Security and privacy Authentication
@@ 1. Introduction
Blockchain technology has been evolving rapidly since its first introduction by Satoshi Nakamoto in 2008 (Nakamoto, 2008). Taking advantage of the decentralized nature of blockchains, Web3 applications have attracted much attention from investors and users (Liu et al., 2022). Decentralized Finance, or DeFi, is a prominent example of Web3 application, and its Total Value Locked (TVL) was around 52 billion USD as of December 2023 (theblock.co, 2024). The total value of the Non-Fungible Token (NFT) marketplaces has exceeded two billion USD (Coinmarketcap, 2024). In January 2024, the total value of assets in the Web3 game Midas Miner was over one billion USD (DappRadar, 2024). The considerable financial activities underscore the enormous potential of the Web3 ecosystem.
To grant the correct permission for end users to access off-chain resources, Web3 applications also need to perform user authentication. Web3 authentication is a challenge-response protocol where the user is identified by a public key (wallet address): the Web3 application sends a specific message to the crypto wallet, which is signed by the user and sent back to the application. The application verifies the signature and authenticates the user if it is valid (Wan et al., 2023). This decentralized approach differs markedly from the traditional authentication protocols of Web2. The most common way to perform Web3 authentication is through MetaMask (met, 2024), a famous Web3 wallet for users to manage their digital assets and perform DeFi transactions. As per statistics, MetaMask had 30 million monthly active users (EarthWeb, 2022). Since the Web3 application is fundamentally a website, in this paper, we use the terms application and website interchangeably.
Due to the popularity of Web3 and the large amount of money involved, the security of Web3 is paramount. Existing research works mainly focus on the security risks associated with Web3 applications. These include vulnerabilities in smart contracts  (Smolka et al., 2023; Sendner et al., 2023; He et al., 2021; Bose et al., 2022; Zhang et al., 2020; Chen et al., 2019; So et al., 2020; Rodler et al., 2019; So et al., 2021; Frank et al., 2020), malicious DeFi attacks  (Babel et al., 2023b; Zhou et al., 2023; McLaughlin et al., 2023; Babel et al., 2023a; Victor and Weintraud, 2021; Zhou et al., 2021; Qin et al., 2021; Daian et al., 2020), and the security of NFT markplaces  (Das et al., 2022; White et al., 2022). The security and anonymity of crypto wallets (Sentana et al., 2023; Yan et al., 2023; Uddin et al., 2021) as well as cryptocurrency scams (Li et al., 2023; He et al., 2023) have also raised concerns. However, to the best of our knowledge, almost no work has paid attention to the security of Web3 authentication.
@@
To detect blind message attacks, we have designed and implemented Web3AuthChecker, a dynamic detection tool that examines the security of the Web3 authentication process. Web3AuthChecker bypasses the web page (front-end) and directly tests the API (back-end). It requests Web3 authentication-related APIs and checks the responses to find vulnerabilities. The tool consists of two modules, 1) Checker, which defines a set of general attack payloads and rules for testing APIs to find potential vulnerabilities, and 2) FlexRequest, which is an HTTP library designed to streamline the testing of identical APIs on different websites. Specifically, to test a website, we load the authentication-related APIs of the website to Checker. Checker then instructs FlexRequest to make a series of requests to these APIs. Finally, Checker examines the data returned by FlexRequest to determine the potential presence of this vulnerability.
To evaluate the impact of Blind Message Attacks, we collected 29 real-world cases of Web3 authentication, including 25 cases of using Web3 authentication for user login and 4 cases for profile update. The 29 cases are from 27 websites, with some encompassing both types. These websites spanning marketplaces, games, and services, sourced from DappRadar and Google searches, reported in January 2024 alone a total transaction volume exceeding 592 million US dollars and over 1.29 million unique active wallets (UAW). By testing the APIs of 29 cases, Web3AuthChecker reports that 75.8% (22/29) of Web3 authentications belonging to 20 websites are vulnerable to blind message attacks. We manually checked all 29 cases and confirmed that the results are accurate. Building upon the foundation of the blind message attack, we have developed more advanced attacks: the Replay Attack renders the session mechanism ineffective, while the Blind Multi-Message Attack allows attackers to acquire user identities across multiple websites in a single attack. In our study of 29 cases, we identified 11 instances of the Replay Attacks and 7 of the Blind Multi-Message Attacks.
@@
To address these vulnerabilities, a complete Web3 authentication update is required for websites and crypto wallets, but this is difficult to implement immediately. Therefore, we proposed a user-side mitigation solution, Web3AuthGuard, to help users immediately mitigate blind message attacks in crypto wallets. When a user logs into a website B, the wallet equipped with Web3AuthGuard will extract a message template from B’s message. Later, when the user attempts to log into a new website A, the wallet will perform a regex match between the message from A and the message template from B. If a match is successful, the wallet will alert the user about the potential blind message attack on that website. We implement Web3AuthGuard in MetaMask’s open-source code. We tested 25 user logins, and Web3AuthGuard successfully addressed 20 of them. Web3AuthGuard is unable to handle the remaining five cases because their vulnerabilities allow an attacker to modify the message body at will.
@@
Following the responsible disclosure policy, we reported our discovered vulnerabilities to the corresponding websites. One of them (LearnBlockchain) has acknowledged and fixed the vulnerability. Additionally, since some vendors did not respond to our security reports, we submitted our findings directly to the CVE program. Currently, two CVE IDs (CVE-2023-50053 and CVE-2023-50059) have been assigned.
@@
  • User-side Mitigation (Web3AuthGuard). We designed and implemented a user-side mitigation solution, Web3AuthGuard, which helps to alert users about a potential Blind Message Attack when the user attempts to log into a new website (Section 7).
  • Open-source Release. To benefit the Web3 community, we will open source our code on GitHub, including the detection tool Web3AuthChecker  [^1] https://github.com/d0scoo1/Web3AuthChecker and a MetaMask-based crypto wallet with built-in Web3AuthGuard  [^2] https://github.com/d0scoo1/Web3AuthGuard.
@@ 2.1. Web3
Web3, also known as Web 3.0, signifies the upcoming evolution of the Internet, moving away from centralized servers and data centers owned by a few large corporations. Instead, it uses a distributed network, that is blockchain to host the data  (Liu et al., 2022). Blockchains are decentralized ledgers that record transactions across many nodes (miners). This ensures that the data cannot be altered without the consensus of the network.
@@
Applications built on the blockchain network are called Web3 applications or decentralized applications (DApps). Web3 applications cover a broad scope, including NFT marketplaces (NFTM), which allow creators and buyers to trade digital assets efficiently and securely; decentralized exchanges (DEX), which enable trading of digital currencies without a central authority; and decentralized finance (DeFi) platforms, which provide financial services such as lending and borrowing. Each Web3 application typically consists of a smart contract and a website. Smart contracts automate operations (transactions) and store critical data on the blockchain, enhancing transparency and security. Websites serve as the user interface and manage off-chain data. While some Web3 applications are embedded within mobile apps like crypto wallets, they fundamentally function as websites. This paper focuses on websites of the Web3 applications and uses the terms application and website interchangeably.
@@
Crypto wallet is a crucial component of the Web3 ecosystem, allowing users to securely store and manage their digital assets (key pairs) using hardware or software (Uddin et al., 2021). Most Web3 applications (websites) are compatible with popular crypto wallets and interact with them through APIs such as Web3.js (web, 2024) or ethers.js (eth, 2023). When a user accesses a website, the website examines the browser for the presence of a Web3 provider, which is the API interface for crypto wallets. If the Web3 provider is detected, the website establishes a connection to it and prompts the user to connect their wallet. After receiving the user approval, the wallet shares the account addresses (public keys) with the website. This facilitates subsequent interactions, including transactions, smart contract operations, and Web3 authentication.
@@ 2.2. Impacts of Unauthorized Access in Web3
Web3 authentication only grants users access to off-chain data, and on-chain transactions usually require additional signatures, so Web3 authentication will have no impact on users’ on-chain assets. However, attackers can still profit from unauthorized access to user accounts. Here, we highlight two significant asset losses: Unlocked Content and Unfair Trading.
Unlocked Content. In NFT marketplaces, some items derive their value from unlocked content. If attackers gain unauthorized access to these accounts, they can directly access the locked content without any cost. Moreover, most Web3 websites do not have anomalous login detection, meaning that the owners may never be notified when their locked content is leaked or stolen.
@@
The key pair in the crypto wallet is randomly generated, so the user’s public key (address) has good anonymity, but an attacker who gains access to a user’s account can link the user’s address to personal information such as email, effectively breaking the anonymity. Moreover, unauthorized access can damage users’ reputations, as attackers may use the user’s account for illegal activities. NFT marketplaces allow users to link their social media accounts, such as Twitter or Facebook. These social media accounts represent important channels for artists to strengthen their connections with users and promote their artworks (White et al., 2022), which attackers can exploit to promote illegal activities, causing reputational damage.
@@ 3. Web3 Authentication
This section first introduces the Web3 authentication process in detail, then specifically focuses on message design and server verification in authentication.
@@ 3.1. Overview
Web3 authentication is primarily utilized in two scenarios: user login and profile update. In the first scenario, the client needs to sign a specific message, which is then verified by the server. Once authenticated, the server will issue a token to the client to sustain a session. In the second scenario, the user signs the updated profile, and the server checks this signature before updating the user’s profile. These two scenarios are independent; some websites may include only one of them, while others incorporate both. The user login is the most common scenario of Web3 authentication and, therefore, this paper mainly focuses on this scenario. Figure 1 illustrates the process of a user logging into a website by Web3 authentication. The process is described below.
@@
The user (front-end) holds the token to access sensitive information of the account (3.1-3.4). To keep the session, the website may store the token in the request header or cookie. This paper only focuses on Web3 authentication, so the details of the session will be omitted.
@@ 3.2. Message Design
In Steps 2.3-2.6, Web3 authentication normally uses the protocol Personal Sign (EIP-191) (Swende and Johnson, 2016). This protocol does not impose any requirement on the format or content of a message, so a Web3 authentication implementation must design the message by itself. We conducted an extensive survey and identified common fields in a message, represented in Augmented Backus–Naur Form (ABNF) as shown in Listing 1. The fields referenced from other RFCs (rfc, 2024) are annotated in the ABNF. These fields can be categorized as variable fields, such as nonce and address, and static fields, like domain and statement. For convenience, the content composed of static fields in the message is referred to as the message body.
In Web3 authentication, each field serves a distinct function, and the design of the message differs across various websites. Listing 2 and Listing 3 provide two examples from well-known NFT marketplaces, opensea.io and foundation.app, respectively. The message from opensea.io is well-designed, including fields such as domain and nonce, whereas the message from foundation.app contains only name and statement, thereby posing security vulnerabilities. Next, we will explain the meaning and function of each field in Listing 1 in detail, using these two examples as references.
@@
The domain field represents the website’s domain name, a crucial field of the message, enabling users to identify the source of message. For example, The message in Listing 2 includes the domain field opensea.io. If a malicious website prompts the user to sign this message, the user can recognize that it comes from opensea.io and refuse to sign it.
@@
The ext field serves as an extension and is not central to this work. Appendix A provides more details.
@@ 3.3. Server Verification
The server verifies the nonce in the message to avoid replay attacks. Depending on the nonce type, the server might need to query the existence of a nonce record (for record-based nonces) or verify whether the timestamp is within an acceptable time range relative to its internal clock (for time-based nonces). Other variable fields on the server such as address will also be verified during this process.
@@
The server verifies the validity of the signature: it uses the user’s address (public key) to decrypt the signature, and then compares the result with the hash of the message. If they match, the signature is deemed valid. Once verified, the server issues an authentication token to authorize user access.

### v4

@@ Abstract.
As the field of Web3 continues its rapid expansion, the security of Web3 authentication, often the gateway to various Web3 applications, becomes increasingly crucial. Despite its widespread use as a login method by numerous Web3 applications, the security risks of Web3 authentication have not received much attention. This paper investigates the vulnerabilities in the Web3 authentication process and proposes a new type of attack, dubbed blind message attacks. In blind message attacks, attackers trick users into blindly signing messages from target applications by exploiting users’ inability to verify the source of messages, thereby achieving unauthorized access to the target application. We have developed Web3AuthChecker, a dynamic detection tool that interacts with Web3 authentication-related APIs to identify vulnerabilities. Our evaluation of real-world Web3 applications shows that a staggering 75.8% (22/29) of Web3 authentication deployments are at risk of blind message attacks. In response to this alarming situation, we implemented Web3AuthGuard on the open-source wallet MetaMask to alert users of potential attacks. Our evaluation results show that Web3AuthGuard can successfully raise alerts in 80% of the tested Web3 authentications. We have responsibly reported our findings to vulnerable websites and have been assigned two CVE IDs.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security; October 14–18, 2024; Salt Lake City, UT, USAbooktitle: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security (CCS ’24), October 14–18, 2024, Salt Lake City, UT, USAdoi: 10.1145/3658644.3670323isbn: 979-8-4007-0636-3/24/10ccs: Security and privacy Authentication
@@ 1. Introduction
Blockchain technology has been evolving rapidly since its first introduction by Satoshi Nakamoto in 2008 (Nakamoto, 2008). Web3 applications, taking advantage of the decentralized nature of blockchain technology, have garnered significant attention from both investors and users (Liu et al., 2022). Decentralized Finance, or DeFi, is a prominent example of the Web3 application, and its Total Value Locked (TVL) was around 52 billion USD as of December 2023 (theblock.co, 2024). The total value of the Non-Fungible Token (NFT) marketplaces has exceeded two billion USD (Coinmarketcap, 2024). In January 2024, the total value of assets in the Web3 game Midas Miner was over one billion USD (DappRadar, 2024). The considerable financial activities underscore the enormous potential of the Web3 ecosystem.
Web3 applications need to perform user authentication to grant the correct permissions for end users to access off-chain resources. Web3 authentication is a challenge-response protocol where the user is identified by a public key (wallet address): the Web3 application sends a specific message to the crypto wallet, signed by the user, and sent back to the application. The application verifies the signature and authenticates the user if it is valid (Wan et al., 2023). This decentralized approach differs markedly from the traditional authentication protocols of Web2. The most common way to perform Web3 authentication is through MetaMask (met, 2024), a famous Web3 wallet for users to manage their digital assets and perform DeFi transactions. As per statistics, MetaMask had 30 million monthly active users (EarthWeb, 2022). Since the Web3 application is fundamentally a website, in this paper, we use the terms application and website interchangeably.
Due to the popularity of Web3 and the large amount of money involved, the security of Web3 is paramount. Existing research mainly focuses on the security risks associated with Web3 applications. These include vulnerabilities in smart contracts  (Smolka et al., 2023; Sendner et al., 2023; He et al., 2021; Bose et al., 2022; Zhang et al., 2020; Chen et al., 2019; So et al., 2020; Rodler et al., 2019; So et al., 2021; Frank et al., 2020), malicious DeFi attacks  (Babel et al., 2023b; Zhou et al., 2023; McLaughlin et al., 2023; Babel et al., 2023a; Victor and Weintraud, 2021; Zhou et al., 2021; Qin et al., 2021; Daian et al., 2020), and the security of NFT marketplaces (Das et al., 2022; White et al., 2022). The security and anonymity of crypto wallets (Sentana et al., 2023; Yan et al., 2023; Uddin et al., 2021), as well as cryptocurrency scams (Li et al., 2023; He et al., 2023) have also raised concerns. However, to the best of our knowledge, almost no work has paid attention to the security of Web3 authentication.
@@
To detect Blind Message Attacks, we have designed and implemented Web3AuthChecker, a dynamic detection tool that examines the security of the Web3 authentication process. Web3AuthChecker bypasses the web page (front-end) and directly tests the API (back-end). It requests Web3 authentication-related APIs and checks the responses to find vulnerabilities. The tool consists of two modules: 1) Checker, which defines a set of general attack payloads and rules for testing APIs to find potential vulnerabilities, and 2) FlexRequest, an HTTP library designed to streamline testing identical APIs on different websites. Specifically, to test a website, we load the authentication-related APIs of the website to Checker. Checker then instructs FlexRequest to make a series of requests to these APIs. Finally, Checker examines the data returned by FlexRequest to identify potential vulnerabilities.
To evaluate the impact of Blind Message Attacks, we collected 29 real-world cases of Web3 authentication, including 25 cases of using Web3 authentication for user login and 4 cases for profile update. The 29 cases are from 27 websites, with some encompassing both types. The websites we studied, which span marketplaces, games, and services, were sourced from DappRadar and Google searches. In January 2024 alone, these sites reported a total transaction volume exceeding 592 million US dollars and over 1.29 million unique active wallets (UAW), underscoring the widespread nature of these vulnerabilities. By testing the APIs of 29 cases, Web3AuthChecker reports that 75.8% (22/29) of Web3 authentications belonging to 20 websites are vulnerable to Blind Message Attacks. We manually checked all 29 cases and confirmed that the results were accurate. Building upon the foundation of the Blind Message Attack, we have developed more advanced attacks: the Replay Attack renders the session mechanism ineffective, while the Blind Multi-Message Attack allows attackers to acquire user identities across multiple websites in a single attack. In our study of 29 cases, we identified 11 instances of the Replay Attacks and 7 of the Blind Multi-Message Attacks.
@@
To address these vulnerabilities, websites and crypto wallets must implement a complete Web3 authentication update, but this is difficult to implement immediately. Therefore, we proposed a user-side mitigation solution, Web3AuthGuard, to help users immediately mitigate Blind Message Attacks in crypto wallets. When a user logs into a website B, the wallet equipped with Web3AuthGuard will extract a message template from B’s message. Later, when the user attempts to log into a new website A, the wallet will perform a regex match between the message from A and the message template from B. If a match is successful, the wallet will alert the user about the potential Blind Message Attack on that website. We implement Web3AuthGuard in MetaMask’s open-source code. We tested 25 user logins, and Web3AuthGuard successfully addressed 20 of them. Web3AuthGuard is unable to handle the remaining five cases because their vulnerabilities allow an attacker to modify the message body at will.
@@
Following the responsible disclosure policy, we reported the vulnerabilities we discovered to the corresponding websites. One of them (LearnBlockchain) has acknowledged and fixed the vulnerability. Additionally, since some vendors did not respond to our security reports, we submitted our findings directly to the CVE program, and two CVE IDs (CVE-2023-50053 and CVE-2023-50059) have been assigned.
@@
  • User-side Mitigation (Web3AuthGuard). We designed and implemented a user-side mitigation solution, Web3AuthGuard, which helps alert users about a potential Blind Message Attack when the user attempts to log into a new website (Section 7).
  • Open-source Release. To benefit the Web3 community, we will open-source our code on GitHub, including the detection tool Web3AuthChecker  [^1] https://github.com/d0scoo1/Web3AuthChecker and a MetaMask-based crypto wallet with built-in Web3AuthGuard  [^2] https://github.com/d0scoo1/Web3AuthGuard.
@@ 2.1. Web3
Web3, also known as Web 3.0, signifies the upcoming evolution of the Internet. It will move away from centralized servers and data centers owned by a few large corporations and instead use a distributed network, blockchain, to host the data  (Liu et al., 2022). Blockchains are decentralized ledgers that record transactions across many nodes (miners). This ensures that the data cannot be altered without the consensus of the network.
@@
Applications built on the blockchain network are called Web3 applications or decentralized applications (DApps). Web3 applications cover a broad scope, including NFT marketplaces (NFTM), which allow creators and buyers to trade digital assets efficiently and securely; decentralized exchanges (DEX), which enable trading of digital currencies without a central authority; and decentralized finance (DeFi) platforms provide financial services such as lending and borrowing. Each Web3 application typically consists of a smart contract and a website. Smart contracts automate operations (transactions) and store critical data on the blockchain, enhancing transparency and security. Websites serve as the user interface and manage off-chain data. While some Web3 applications are embedded within mobile apps like crypto wallets, they fundamentally function as websites. This paper focuses on websites of Web3 applications and uses the terms application and website interchangeably.
@@
Crypto wallets are a crucial component of the Web3 ecosystem, allowing users to securely store and manage their digital assets (key pairs) using hardware or software (Uddin et al., 2021). Most Web3 applications (websites) are compatible with popular crypto wallets and interact with them through APIs such as Web3.js (web, 2024) or ethers.js (eth, 2023). When a user accesses a website, the website examines the browser for the presence of a Web3 provider, which is the API interface for crypto wallets. If the Web3 provider is detected, the website establishes a connection to it and prompts the user to connect their wallet. After user approval, the wallet shares the account addresses (public keys) with the website. This facilitates subsequent interactions, including transactions, smart contract operations, and Web3 authentication.
@@ 2.2. Impacts of Unauthorized Access in Web3
Web3 authentication only grants users access to off-chain data, and on-chain transactions usually require additional signatures so that Web3 authentication will not impact users’ on-chain assets. However, attackers can still profit from unauthorized access to user accounts. Here, we highlight two significant asset losses: Unlocked Content and Unfair Trading.
Unlocked Content. In NFT marketplaces, some items are valued from unlocked content. If attackers gain unauthorized access to these accounts, they can directly access the locked content without any cost. Moreover, most Web3 websites do not have abnormal login detection, meaning that the owners may never be notified when their locked content is leaked or stolen.
@@
The key pair in the crypto wallet is randomly generated, so the user’s public key (address) has good anonymity, but an attacker who gains access to a user’s account can link the user’s address to personal information such as email, effectively breaking the anonymity. Moreover, unauthorized access can damage users’ reputations, as attackers may use the user’s account for illegal activities. NFT marketplaces allow users to link their social media accounts, such as Twitter or Facebook. These social media accounts represent important channels for artists to strengthen their connections with users and promote their artworks (White et al., 2022), which attackers can exploit to encourage illegal activities, causing reputational damage.
@@ 3. Web3 Authentication
This section briefly introduces the Web3 authentication process, then specifically focuses on message design and server verification in authentication.
@@ 3.1. Overview
Web3 authentication is primarily utilized in two scenarios: user login and profile update. In the first scenario, the client needs to sign a specific message, which is then verified by the server. Once authenticated, the server will issue a token to the client to sustain a session. In the second scenario, the user signs the updated profile, and the server checks this signature before updating the user’s profile. These two scenarios are independent; some websites may include only one of them, while others incorporate both. The user login is the most common scenario of Web3 authentication, and therefore, this paper mainly focuses on this scenario. Figure 1 illustrates the process of a user logging into a website by Web3 authentication. The process is described below.
@@
The user (front-end) holds the token to access protected information (3.1-3.4). To keep the session, the website may store the token in the request header or cookie. This paper only focuses on Web3 authentication, so the session details are omitted.
@@ 3.2. Message Design
In Steps 2.3-2.6, Web3 authentication normally uses the protocol Personal Sign (EIP-191) (Swende and Johnson, 2016). This protocol does not impose any requirement on the format or content of a message, so Web3 applications must implement their own messages for authentication. We conducted an extensive survey and identified common fields in messages, represented in Augmented Backus–Naur Form (ABNF) as shown in Listing 1. The fields referenced from other RFCs (rfc, 2024) are annotated in the ABNF. These fields can be categorized as variable fields, such as nonce and address, and static fields, like domain and statement. For convenience, the content composed of static fields in the message is referred to as the message body.
In Web3 authentication, each field serves a distinct function, and the design of the message differs across various websites. Listing 2 and Listing 3 provide two examples from well-known NFT marketplaces, opensea.io and foundation.app, respectively. The message from opensea.io is well-designed, including fields such as domain and nonce. In contrast, the message from foundation.app contains only name and statement, thereby posing security vulnerabilities. Next, we will explain the meaning and function of each field in Listing 1 in detail, using these two examples as references.
@@
The domain field is a critical element of the message, representing the website’s domain name. It plays a pivotal role in message security, enabling users to verify the source of the message. For example, the message in Listing 2 includes the domain field opensea.io. If a malicious website prompts the user to sign this message, the user can quickly recognize that it comes from opensea.io and refuse to sign it.
@@
The ext field serves as an extension and is not central to this work. For more details, please refer to EIP-4361 (Chang et al., 2021).
@@ 3.3. Server Verification
The server verifies the nonce in the message to avoid replay attacks. Depending on the nonce type, the server might need to query the existence of a nonce record (for record-based nonces) or verify whether the timestamp is within an acceptable time range relative to its internal clock (for time-based nonces). Other variable fields on the server, such as address, will also be verified during this process.
@@
The server verifies the validity of the signature: it uses the user’s address (public key) to decrypt the signature. Then, it compares the result with the hash of the message. If they match, the signature is deemed valid. Once verified, the server issues an authentication token to authorize user access.